OpenStudy (mazehaq):

Given sinθ = -2/3, π, < θ < 3π/2, find the values of the remaining trigonometric functions of θ. Help please! We didn't go over this in class & this is all foreign for me.

OpenStudy (anonymous):

Now solve by ur selves

OpenStudy (mazehaq):

How did you get -sqrt (5)? I don't understand the concept, are we trying to find reference angles?

OpenStudy (anonymous):

\[\sin(\theta)=\frac{ opposite }{ hypotenus }=\frac{ -2 }{ 3 }\]

OpenStudy (anonymous):

Use Pythagoras theorem

OpenStudy (mazehaq):

OH! I kinda get it! So it's like SOH CAH TOA, right? & we're looking for CAH & TOA?

OpenStudy (mazehaq):

The theorem is x^2 + y^2 = 1 or the A^2 + B^2 = C^2, I'm mixing up some stuff.

OpenStudy (anonymous):

\[(-2)^{2}+m^{2}=3^{2}...thus...m=\sqrt{5}\] But m is -ve since theta is third quadrant angle.

OpenStudy (anonymous):

Now can u do that?

OpenStudy (mazehaq):

I get it better!